use the pythagorean theorem to find the indicated quantity. Round your answer to the nearest hundredth.
c^2 = 5^2 + 19^2

Answer by checkley77(12844)

c^2=5^2+19^2
c^2=25+361
c^2=386
c=sqrt386
c=19.647 ans.